# Project page: https://github.com/zhang-tao-whu/PCM

### v3

@@ 5 Conclusion
This paper introduces a Mamba-based point cloud network named Point Cloud Mamba, which, for the first time, outperforms the SOTA point-based method PointNeXt. Point Cloud Mamba incorporates several novel techniques to help Mamba better model point cloud data. Firstly, we propose Consistent Traverse Serialization to convert 3D point cloud data into 1D point sequences that Mamba can handle, ensuring that neighboring points in the sequence are also spatially adjacent. Secondly, we aid Mamba in handling point sequences serialized in different orders by introducing order prompts containing sequence arrangement rules. Finally, we propose a simple yet effective positional encoding based on spatial coordinate mapping. Our Point Cloud Mamba achieves SOTA performance on the ScanObjectNN, ModelNet40, ShapeNetPart, and S3DIS datasets. Adding a stronger local feature extractor, our method also outperforms previous STOA methods by a large margin (over 8% mIoU improvements) on the S3DIS dataset.
Limitations and future work directions. PCM successfully introduces Mamba into point cloud analysis and surpasses modern point-based methods like PointNeXt. However, there are still some limitations that need to be addressed. For large-scale point clouds (i.e., $\geq$ 100k points), such as in the S3DIS dataset, Mamba struggles to handle such long sequences during training due to the scan-based computational approach used to accelerate training. Therefore, it is necessary to crop the input point clouds, but this introduces a gap between training and inference for global modeling architecture PCM. In addition, how to better combine local feature extractors with PCM is also worth trying. Moreover, there are still several directions to explore when adopting PCM in out-door point cloud scene, where the point inputs are huge and more complex. We will focus on addressing these challenges in our future work.
Appendix 0.A More Experiment Results.
Comparison with naive mamba-based architecture. We compare our proposed PCM with the naive Mamba-based architecture on ScanObjectNN [52], ModelNet40 [63], and ShapeNetPart [72] to validate the effectiveness of our design. The comparison results on ScanObjectNN are shown in Tab. 11. It is worth noting that even our proposed naive Mamba-based architecture outperformed the contemporaneous work PointMamba [28] 2.7 OA due to the combination of local modeling and global modeling. However, thanks to our proposed consistent traverse serialization strategy, order prompt, simple positional embedding, and more reasonable architecture settings, our PCM still surpasses the naive architecture with 2.8 OA and 2.9 mAcc on ScanObjectNN. Moreover, the accuracy in almost all categories is higher than in the naive Mamba-based architecture. The comparison results on ModelNet40 are shown in Tab. 12. PCM also outperforms the naive Mamba-based architecture with 1.5 OA and 2.4 mAcc.
The comparison of part segmentation performance on the ShapeNetPart dataset is shown in Tab. 13. The naive Mamba-based architecture surpasses the contemporaneous work PointMamba with 0.8 Ins. mIoU and 0.9 Cls. mIoU. PCM also exceeds the naive architecture with 0.4 Ins. mIoU and 0.5 Cls. mIoU.
Comparison of the multiple serialization strategy with the single serialization strategy. In the main paper, we compared the impact of the multiple serialization strategy and single serialization strategy on point cloud classification performance. The multiple serialization strategy yielded improvements of 0.42 OA and 0.87 mAcc compared to the single serialization strategy. As shown in Tab. 14, on ShapeNetPart, the multiple serialization strategy also led to performance enhancements of 0.4 Ins. mIoU and 0.7 Cls. mIoU in part segmentation.
Comparison of the Parameters, FLOPs, and Throughput. We summarize our proposed PCM’s parameter, computational complexity, and throughput, as shown in Tab. 15. Our proposed PCM-Tiny outperformed PointMLP 1.5 OA and 1.1 mAcc with only 52% parameters and 35% computational complexity. PCM surpassed PointMLP [35] 2.7 OA and 2.7 mAcc with a larger amount of parameters and computational complexity. However, due to multiple reorderings, the throughput of PCM is not advantageous; even though PCM-Tiny has fewer parameters than PointMLP, its throughput is still lower than PointMLP.
Figure 6: The failure cases of PCM. Incorrect areas are highlighted by red rectangles.
Figure 7: More visualization results.
Appendix 0.B Visualization Results
Failure cases. Fig. 6 illustrates some instances of PCM failure. For example, on the left side, PCM performs poorly for certain smaller object parts, often misclassifying them as parts of larger, similar ones. PCM is also susceptible to issues when the point cloud has numerous missing points, as evidenced by the car in the middle.
More visualization results. In Fig. 7, we present more visualization results. Even when dealing with elongated or flattened objects, PCM still achieves good results. This demonstrates that serializing the point cloud into a point sequence and then using Mamba to model global features is feasible and effective, even when the point cloud is distributed irregularly in space.
Appendix 0.C Further works.
In future work, we will focus on how to utilize Mamba for the global modeling of large-scale point cloud scenes. Since Mamba employs scan-based computation during training to enhance parallelism, which incurs quadratic computational complexity, it is not feasible to directly process the whole point cloud with Mamba during training. However, during testing, the entire point cloud is often processed at once, creating a substantial gap between training and testing and thereby limiting the performance of Mamba-based methods. We will explore strategies to bridge this gap, such as through scalable serialization methods.

### v4

@@ Conclusion
This paper introduces a Mamba-based point cloud network named Point Cloud Mamba, which, for the first time, outperforms the SOTA point-based method PointNeXt and transformer-based method PTv3. Point Cloud Mamba incorporates several novel techniques to help Mamba better model point cloud data. Firstly, we propose Consistent Traverse Serialization to convert 3D point cloud data into 1D point sequences that Mamba can handle, ensuring that neighboring points in the sequence are also spatially adjacent. Secondly, we aid Mamba in handling point sequences serialized in different orders by introducing order prompts containing sequence arrangement rules. Finally, we propose a simple yet effective positional encoding based on spatial coordinate mapping. Our Point Cloud Mamba achieves SOTA performance on the ScanObjectNN, ModelNet40, ShapeNetPart, and S3DIS datasets. Adding a stronger local feature extractor, our method also outperforms previous STOA methods by a large margin on the S3DIS dataset.